the no. of evan divisors of 1600 are:


(a)18 (b)21 (c)3 (d)none


ans: a

Grade:12

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear shefali

1600=2752

 

so number of all divisors =(7+1)(2+1) =21

 but these diviser also include    1,50,52 ie  1,5,25

  which is not even 

so even divisior =21-3 =18

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

Anjana Priya
5 years ago
I think the process mentioned above is correct, but the calculation is not clear  That is,
      1600=2^6 × 5^2           (not 2 raised by 7)
Number of divisors=(6+1)(2+1)=21
And finally minus 3 as explained above.
So, the final answer is 21-3=18
 
Thanks  for considering my view. 
Regards.